# DeepFRC: An End-to-End Deep Learning Model for Functional Registration and Classification

### v2

@@ 4 Experiments
We present the results of functional data registration and classification using DeepFRC, evaluated on both simulated and real-world datasets. Simulated data provides insights into registration, reconstruction, and classification progression during training. Real-world datasets allow for performance comparisons with state-of-the-art methods. To test model generalization, training sets are deliberately smaller than test sets.
@@
We evaluate registration performance using the metric $\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})$ within the EFDA framework. This metric is applicable only to simulated datasets, as they provide true warping functions $\bm{\gamma}^{*}$ or underlying true curves. Another metric for alignment assessment is the adjusted total variance (ATV) defined as: $\text{ATV}=\frac{1}{\binom{C}{2}}\sum_{1\leq u<v\leq C}\frac{\text{TV}_{uv}}{d(\text{mean}(\tilde{x}_{u}),\text{mean}(\tilde{x}_{v}))},$ where $\text{TV}_{uv}$ (defined in (Chen and Srivastava, 2021)) is the total variation between classes $u$ and $v$, and $\text{mean}(\tilde{x})$ denotes the mean function of the aligned curve $\tilde{x}$. ATV accounts for the distance between class mean functions, making it more suitable for alignment evaluation of similar classes. Lower values of $\Delta Q_{\text{reg}}$ and ATV indicate better alignment. To assess DeepFRC’s ability to recover the underlying smooth process of aligned functional data, we compute the Pearson’s correlation coefficient $\rho(\bm{c}^{*},\hat{\bm{c}})$ between the true and estimated Fourier basis coefficients. This metric is used only in simulations, as $\bm{c}^{*}$ can only be computed when the true curves are known. Higher values of $\rho$ indicate better reconstruction. For classification performance, we evaluate accuracy (ACC) and $F_{1}$-score with the latter provides a more balanced measure of performance. We use the macro-averaged $F_{1}$-score (Sokolova and Lapalme, 2009) for multi-class classification. Higher values of ACC and $F_{1}$-score indicate better classification performance.
@@ 4.1 Simulation
We generate a balanced two-class functional dataset $\{x_{i}(t),y_{i}\}_{i=1}^{N}$, where each function $x_{i}(t)$ combines amplitude and phase variation. Amplitude is modeled by sums of Gaussian bumps, while phase variation is introduced via nonlinear warping functions $\gamma_{i}(t)$. Class-specific parameters and full generation details are provided in Appendix E.1. We simulate $N=6000$ functions, each evaluated at 1000 time points.
@@ Results.
Figure 2 illustrates the effect of iterative optimization in DeepFRC on the simulated dataset. Figure 2(a) shows the raw curves of two classes (yellow and blue) with registration issues, which are challenging to distinguish visually. As shown in Figures 2(a)-(d), alignment improves progressively with the number of epochs, evidenced by the curves becoming more centered for each class, making them easier to differentiate. This process visually shows the mutual benefit of alignment and classification. Figure 2(e) demonstrates that the Pearson correlation coefficient $\rho(\bm{c}^{*},\hat{\bm{c}})$ increases rapidly over epochs, reaching nearly 1 by epoch 50, indicating DeepFRC’s strong reconstruction capability for the true unwarped functional data. Figures 2(f)-(g) show the progression of alignment and classification performance, all improving steadily with increasing epochs. Finally, Figure 2(h) presents the total loss function progression, which includes the registration loss (Eq. (10)) and classification loss (Eq. (11)). All three loss components decrease sharply before epoch 50 and stabilize near zero thereafter. Visualization of alignment and numerical results on the simulated test set are provided in Figure A1 of Appendix E.1, demonstrating excellent generalization performance.
@@ 4.2 Real Data Application
We evaluate our method on four publicly available datasets from the UCR Time Series Classification Archive, selected for their phase variation characteristics and common use in functional data analysis (Ramsay and Silverman, 2002, 2005): (1) Wave (motion): The Wave dataset contains 8 gesture classes recorded by accelerometers along three axes. We use the X-axis recordings from classes 2 and 8, yielding 1120 curves with 315 time points each. (2) Yoga (image-derived): The Yoga dataset features 1D time-series representations of yoga poses, derived by measuring the distance from the silhouette to the image center. It contains 3300 samples with 426 time points each. (3) Symbol (handwriting): The Symbol dataset records pen trajectories as participants replicate symbols. We use two subsets: a binary-class subset with 343 samples, and a three-class subset with 510 samples. Each sequence has 398 time points. The Wave and Symbol datasets are class-balanced, while Yoga is imbalanced. Further details are provided in Appendix E.2.
@@
We benchmark DeepFRC against several deep learning models for functional data analysis, including a joint registration-classification model (TTN (Lohit et al., 2019)), a registration-only model (SrvfRegNet (Chen and Srivastava, 2021)), and four classification models: $\text{FCNN}_{\text{raw}}$, $\text{FCNN}_{\text{fourier}}$, FuncNN (Thind et al., 2020), and ADAFNN (Yao et al., 2021). To ensure a fair comparison, we evaluate sequential registration-classification pipelines by combining SrvfRegNet with each classification model (excluding TTN, which is already joint). A detailed description of all baselines appears in Appendix E.3.
@@
All models are trained to optimal performance; implementation details are given in Appendix E.4. Table 1 reports both alignment and classification metrics across four real-world datasets. Among all models, only DeepFRC and TTN jointly perform registration and classification. DeepFRC outperforms TTN by using elastic FDA for enhanced functional smoothness, delivering superior registration across all datasets and significant classification improvements on Symbol data. Unlike pipelines combining SrvfRegNet and classification models - which disregard label information during registration - DeepFRC achieves superior alignment, leading to more interpretable transformations and improved downstream classification accuracy. Figure A2 (Appendix E.4) plots the training loss versus epochs for DeepFRC across all datasets, confirming convergence.
Table 1: Quantitative comparison of registration and classification performance with state-of-the-art approaches across four real datasets. Bold indicates best results.
Wave Yoga Symbol (2 classes) Symbol (3 classes)
Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963
TTN 6.3 94.7% 0.948 57.7 89.4% 0.904 8.6 92.0% 0.918 4.5 93.3% 0.933
SrvfRegNet+$\text{FCNN}_{\text{raw}}$ 7.3 94.6% 0.947 136.0 81.0% 0.830 14.8 94.5% 0.942 6.5 94.7% 0.947
SrvfRegNet+$\text{FCNN}_{\text{fourier}}$ 7.3 94.9% 0.950 136.0 84.0% 0.852 14.8 95.0% 0.949 6.5 96.0% 0.959
SrvfRegNet+FuncNN 7.3 95.7% 0.957 136.0 89.0% 0.908 14.8 93.5% 0.933 6.5 94.7% 0.947
SrvfRegNet+ADAFNN 7.3 94.6% 0.949 136.0 73.4% 0.753 14.8 89.0% 0.894 6.5 94.3% 0.943
@@
Figure 3 compares alignment quality on the Symbol (3 classes) dataset. DeepFRC produces smooth, class-separated functional alignments, while TTN distorts class-specific trajectories (e.g., purple), and SrvfRegNet loses inter-class separation (e.g., blue vs. yellow). Both baselines significantly degrade interpretability (see also Figure A3, Appendix E.4). DeepFRC’s superior alignment improves inference reliability and preserves generalization performance - critical for real-world applications.
@@
Table 2 presents ablation study results that separately evaluate the registration and classification components of our method. Removing the classification module (DeepFRC w/o Class.) degrades registration accuracy, while removing registration (DeepFRC w/o Reg.) harms classification - especially on Yoga and Symbol (2 classes). These results highlight the synergy between registration and classification within DeepFRC’s unified architecture.
Table 2: Ablation study: contributions of registration and classification components.
Wave Yoga Symbol (2 classes) Symbol (3 classes)
Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963
DeepFRC w/o Class. 7.3 136.0 14.8 6.5
DeepFRC w/o Reg. 94.4% 0.946 83.1% 0.846 91.0% 0.905 94.7% 0.947
@@
We assess DeepFRC’s robustness to the choice of basis family and size $K$ across four real datasets in Figure 4. The left panel shows performance is stable across basis types (Fourier, B-spline, polynomial). We attribute this robustness to two factors: (a) joint training adapts alignment to the inductive biases of each basis, and (b) class-aware registration loss guides the model toward discriminative structures, reducing basis-specific limitations. The right panel shows that increasing $K$ improves performance up to around $K=100$, after which it plateaus. We believe this invariance arises from (a) upstream registration reducing functional variability, and (b) downstream layers adaptively re-weighting or pruning redundant basis functions.
@@ 5 Discussion
Robustness to Missing Data.
We evaluated DeepFRC’s tolerance to missing observations by randomly removing 5-10% of data points from half of each real-world dataset (Section 4.2), imputing gaps via Fourier splines (Wahba, 1990). Under identical hyperparameters, DeepFRC achieved comparable registration and classification performance to complete data (Figure A4, Appendix E.4 ), except on the noisy Yoga dataset, where alignment degraded slightly at 10% missingness. Results confirm robustness to minor missing data with proper smoothing.
@@ 6 Conclusion
This work introduces DeepFRC, a robust framework for joint functional registration and classification, supported by theoretical guarantees and extensive empirical validation. DeepFRC outperforms state-of-the-art methods and remains stable across real-world scenarios, including missing data and both small- and large-scale datasets. Assuming a shared hidden warping pattern across dimensions, DeepFRC naturally extends to $d$-dimensional functional data ($d\geq 2$) by replacing the input $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$ with $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n\times d}$, without introducing additional assumptions.
Broader Impacts and Limitations.
This is a methodological and theoretical contribution without direct societal implications. While DeepFRC exhibits empirical robustness to the choice and dimensionality of basis functions, a theoretical formalization of this invariance remains an open direction. Additionally, its performance may decline on highly volatile functional data, where the current objective is sensitive to rapid directional fluctuations. This suggests the need for more adaptive mechanisms or robust loss formulations to better accommodate extreme signal variability. Future work could incorporate signal complexity awareness into the model architecture or objective design to further improve generalization.
@@ Appendix B Model Selection of DeepFRC
For $10<\delta_{3}<$ 1, there exists a $T>0$ such that:
@@
$\left|Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)-Q_{i}(\hat{\gamma}_{i})\left(\frac{k}{T}\right)\right|\leq M_{-1}^{(3)}\varepsilon,\quad\text{where }M_{-1}^{(3)}\text{ is a constant.}$.
@@
The loss function is obviously continuous for $\psi_{ij}\geq\epsilon_{0}$, as $\frac{1}{z}$ is continuous everywhere except at zero. Similar to Eqs. (A.1) and (A.2), the gradient of $l$ with respect to any $\theta\in\Theta$ can be easily derived. These gradients can be expressed as products of gradients from ReLU, max pooling, etc., which are Lipschitz continuous. Therefore, the theorem follows directly from Theorem 3.8 in Hardt et al. [2016].
@@ Appendix E Experimental Details
E.1 Description of Simulated Datasets
@@
where the parameters $\{a_{i}^{(cj)},\mu^{(cj)},\sigma_{i}^{(cj)}\}_{j=1}^{2}$ for $c=0,1$ are summarized in in the following table. We split the balanced dataset into training, validation, and test sets with sizes 1600, 400, and 4000, respectively.
@@
$\sigma_{i}$ $0.06+U(-0.003,0.003)$ $0.075+U(-0.003,0.003)$ $0.06+U(-0.003,0.003)$ $0.1+U(-0.003,0.003)$
[…]
E.2 Description of Real Datasets
The Four time series datasets from the UCR Time Series Classification Archive (https://www.timeseriesclassification.com/index.php):
@@ E.2 Description of Real Datasets
E.3 Introduction of Baselines

### v3

@@ 4 Experiments
We present the results of functional data registration and classification using DeepFRC, evaluated on both simulated and real-world datasets. Simulated data provides insights into registration, reconstruction, and classification progression during training. Real-world datasets allow for performance comparisons with state-of-the-art methods.
@@
We evaluate our method using metrics for registration, reconstruction, and classification. Registration quality is assessed by the alignment error ($\Delta Q_{\text{reg}}$) between estimated and true warping functions on simulated data, and by the Adjusted Total Variance (ATV) (Chen and Srivastava, 2021) on both simulated and real data, where lower values indicate better alignment. The fidelity of the recovered smooth process is measured via the correlation between true and estimated basis coefficients (simulated data only). Classification performance is reported using accuracy (ACC) and the macro-averaged $F_{1}$-score (Sokolova and Lapalme, 2009), with higher values being better. Detailed metric definitions are provided in Appendix E.1.
@@ 4.1 Simulation
We generate a balanced two-class functional dataset $\{x_{i}(t),y_{i}\}_{i=1}^{N}$, where each function $x_{i}(t)$ combines amplitude and phase variation. Amplitude is modeled by sums of Gaussian bumps, while phase variation is introduced via nonlinear warping functions $\gamma_{i}(t)$. Class-specific parameters and full generation details are provided in Appendix E.2.
@@ Results.
[…]
@@ 4.2 Real Data Application
We evaluate our method on five publicly available datasets, selected for their prevalence in functional data analysis (Ramsay and Silverman, 2002, 2005) and relevance to phase variation: Wave, Yoga, Symbol, and MotionSense (Malekzadeh et al., 2019). Wave, Yoga, and Symbol are sourced from the UCR Time Series Classification Archive, while MotionSense is from Kaggle. For the Symbol dataset, we use both binary and three-class subsets. The Wave, Symbol, and MotionSense datasets are class-balanced, whereas Yoga is imbalanced. In terms of dimensionality, Wave, Yoga, and Symbol are one-dimensional, and MotionSense is three-dimensional. Further details are provided in Appendix E.3.
@@
We benchmark DeepFRC against several deep learning models for functional data analysis, including a joint registration-classification model (TTN (Lohit et al., 2019)), a registration-only model (SrvfRegNet (Chen and Srivastava, 2021)), and five classification models: $\text{FCNN}_{\text{raw}}$, $\text{FCNN}_{\text{fourier}}$, FuncNN (Thind et al., 2020), ADAFNN (Yao et al., 2021), and TSLANet (Eldele et al., 2024). To ensure a fair comparison, we evaluate sequential registration-classification pipelines by combining SrvfRegNet with each classification model (excluding TTN, which is already joint). A detailed description of all baseline methods and implementation details are provided in Appendices E.4 and E.5, respectively.
@@
Table 1 reports both alignment and classification metrics across five real-world datasets. Among all models, only DeepFRC and TTN jointly perform registration and classification. DeepFRC outperforms TTN by using elastic FDA for enhanced functional smoothness, delivering superior registration across all datasets and significant classification improvements on Symbol and MotionSense data. Unlike pipelines combining SrvfRegNet and classification models - which disregard label information during registration - DeepFRC achieves superior alignment, leading to more interpretable transformations and improved downstream classification accuracy. We observe that DeepFRC achieves classification accuracy comparable to TSLANet, a recent state-of-the-art time-series model that outperforms many transformer-based alternatives, across all datasets. Figure A2 (Appendix E.5) plots the training loss versus epochs for DeepFRC across all datasets, confirming convergence.
Table 1: Quantitative comparison of registration and classification performance with state-of-the-art approaches across five real datasets. Bold indicates best results.
Wave Yoga Symbol (2 classes) Symbol (3 classes) MotionSense
Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963 25.0 95.0% 0.952
TTN 6.3 94.7% 0.948 57.7 89.4% 0.904 8.6 92.0% 0.918 4.5 93.3% 0.933 35.1 85.0% 0.857
[…]
SrvfRegNet+FuncNN 7.3 95.7% 0.957 136.0 89.0% 0.908 14.8 93.5% 0.933 6.5 94.7% 0.947 37.7 90.0% 0.889
SrvfRegNet+ADAFNN 7.3 94.6% 0.949 136.0 73.4% 0.753 14.8 89.0% 0.894 6.5 94.3% 0.943 37.7 85.0% 0.857
SrvfRegNet+TSLANet 7.3 96.4% 0.961 136.0 89.3% 0.884 14.8 95.5% 0.955 6.5 96.3% 0.960 37.7 95.0% 0.952
@@
Figure 3 compares alignment quality on the Symbol (3 classes) dataset. DeepFRC produces smooth, class-separated functional alignments, while TTN distorts class-specific trajectories (e.g., purple), and SrvfRegNet loses inter-class separation (e.g., blue vs. yellow). Both baselines significantly degrade interpretability (see also Figures A3 and A4, Appendix E.5). DeepFRC’s superior alignment improves inference reliability and preserves generalization performance - critical for real-world applications.
@@
The ablation study in Table 2 evaluates the individual contributions of the core components of our method: the Neural Deformation Operator (N.D.O.) for registration, the Spectral Representation (S.R.), and the Classifier Network (C.N.). The results demonstrate a clear synergy between these modules. Removing the registration component (DeepFRC w/o N.D.O.) significantly degrades classification performance, particularly on the Yoga, Symbol (2-class), and MotionSense datasets. Conversely, removing the classifier (DeepFRC w/o C.N.) impairs registration accuracy. The removal of the spectral representation (DeepFRC w/o S.R.) adversely affects both tasks across all datasets. These findings underscore the importance of the spectral representation and the tight coupling between registration and classification within DeepFRC’s unified architecture.
Table 2: Ablation study: contributions of three components in DeepFRC.
Wave Yoga Symbol (2 classes) Symbol (3 classes) MotionSense
Model ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score ATV ACC $F_{1}$-score
DeepFRC 5.6 96.4% 0.965 16.2 89.8% 0.909 4.8 96.0% 0.959 3.2 96.3% 0.963 25.0 95.0% 0.952
DeepFRC w/o N.D.O. 94.4% 0.946 83.1% 0.846 91.0% 0.905 94.7% 0.947 90.0% 0.909
DeepFRC w/o S.R. 5.8 95.3% 0.955 17.7 89.2% 0.903 5.3 94.5% 0.945 3.3 93.3% 0.933 28.8 90.0% 0.900
DeepFRC w/o C.N. 7.3 136.0 14.8 6.5 37.7
@@
We assess DeepFRC’s robustness to the choice of basis family and size $K$ across five real datasets in Figure 4. The left panel shows performance is stable across basis types (Fourier, B-spline, polynomial). We attribute this robustness to two factors: (a) joint training adapts alignment to the inductive biases of each basis, and (b) class-aware registration loss guides the model toward discriminative structures, reducing basis-specific limitations. The right panel shows that increasing $K$ improves performance up to around $K=100$, after which it plateaus. We believe this invariance arises from (a) upstream registration reducing functional variability, and (b) downstream layers adaptively re-weighting or pruning redundant basis functions.
@@ 5 Discussion
Computational Efficiency.
DeepFRC offers significant computational advantages during inference, with an average runtime of 60s on our real-world datasets—dramatically faster than the traditional registration method DTW (1000s) and competitive with deep baselines like TTN (15s) and SrvfRegNet hybrids (90–100s). This efficiency is achieved by eschewing pairwise dynamic programming. Although training time is on par with other deep models, DeepFRC achieves superior accuracy, striking an excellent trade-off between performance and computational cost.
Robustness to Noise and Reconstruction Accuracy.
We evaluate robustness by injecting Gaussian noise ($\sigma=0,0.05,0.10,0.20$) into synthetic data and measuring reconstruction quality via the Pearson correlation between true and estimated basis coefficients. As shown in Table A3, DeepFRC maintains high correlation under low-to-moderate noise, outperforming TTN and exhibiting greater stability than SrvfRegNet. Even under high noise ($\sigma=0.20$), its performance degrades gracefully while surpassing competitors, indicating that the smooth basis representation and joint optimization confer strong resilience.
@@
Robustness to Missing Data.
We evaluated DeepFRC’s tolerance to missing observations by randomly removing 5-10% of data points from half of each real-world dataset (Section 4.2), imputing gaps via Fourier splines (Wahba, 1990). Under identical hyperparameters, DeepFRC achieved comparable registration and classification performance to complete data (Figure A5, Appendix E.5 ).
@@ 6 Conclusion
We introduced DeepFRC, a unified framework for joint functional registration and classification with theoretical guarantees and comprehensive empirical validation. DeepFRC consistently outperforms state-of-the-art baselines and remains robust across real-world and simulated settings with noise, missing values, and scale variation. While empirically invariant to the choice of basis functions, its theoretical underpinnings remain open. A key limitation is reduced performance on highly volatile trajectories, where the objective is sensitive to rapid directional changes. Future work will focus on adaptive architectures and robust loss designs that explicitly capture signal complexity to enhance generalization.
@@ Appendix B Model Selection of DeepFRC
For $0<\delta_{3}<1$, there exists a $T>0$ such that:
@@
$\left|Q_{i}(\gamma_{i})\left(\frac{k}{T}\right)-Q_{i}(\hat{\gamma}_{i})\left(\frac{k}{T}\right)\right|\leq M_{-1}^{(3)}\varepsilon,\quad\text{where }M_{-1}^{(3)}\text{ is a constant.}$
@@
The loss function is obviously continuous for $\psi_{ij}\geq\epsilon_{0}$, as $\frac{1}{z}$ is continuous everywhere except at zero. Similar to Eqs. (A.1) and (A.2), the gradient of $l$ with respect to any $\theta\in\Theta$ can be easily derived. These gradients can be expressed as products of gradients from ReLU, max pooling, etc., which are Lipschitz continuous. We note that both the 1D linear interpolation and a Fourier basis expansion in the stage of spectral representation in DeepFRC satisfy the Lipschitz continuity condition. Therefore, the theorem follows directly from Theorem 3.8 in Hardt et al. [2016].
@@ Appendix E Experimental Details
E.1 Evaluation Metrics
Registration Metrics:
  • Registration Error ($\Delta Q_{\text{reg}}$): This metric, defined within the EFDA framework, quantifies the discrepancy between the estimated warping $\hat{\bm{\gamma}}$ and the ground-truth $\bm{\gamma}^{*}$. It is applicable only to simulated datasets where $\bm{\gamma}^{*}$ is known. Lower values indicate better alignment.
[…]
Reconstruction Metric:
[…]
Classification Metrics:
  • Accuracy (ACC): The proportion of correctly classified samples.
  • Macro-averaged $F_{1}$-score: Provides a balanced measure of performance across classes, especially important for imbalanced datasets [Sokolova and Lapalme, 2009].
Higher values for both ACC and $F_{1}$-score indicate superior classification performance.
E.2 Description of Simulated Datasets
@@
where the parameters $\{a_{i}^{(cj)},\mu^{(cj)},\sigma_{i}^{(cj)}\}_{j=1}^{2}$ for $c=0,1$ are summarized in in the following table. We simulate $N=6000$ functions, each evaluated at 1000 time points. We split the balanced dataset into training, validation, and test sets with sizes 1600, 400, and 4000, respectively.
@@
$\sigma_{i}$ 0.06+U(0.003,0.003)0.06+U(-0.003,0.003) 0.075+U(0.003,0.003)0.075+U(-0.003,0.003) 0.06+U(0.003,0.003)0.06+U(-0.003,0.003) 0.1+U(0.003,0.003)0.1+U(-0.003,0.003)
Figure A1: Visualization of alignment by DeepFRC on simulated test data, with evaluation metrics $\rho=0.976$, $\Delta Q_{\text{reg}}=0.056$, ATV=0.504, $\text{ACC}=1.000$ and $F_{1}\text{-score}=1.000$.
E.3 Description of Real Datasets
The first four time series datasets are one-dimensional functional data, from the UCR Time Series Classification Archive (www.timeseriesclassification.com), and the last dataset is three-dimensional functional data, from Kaggle (www.kaggle.com):
@@ E.3 Description of Real Datasets
The MotionSense [Malekzadeh et al., 2019] dataset involves multivariate time-series signals recorded from smartphone sensors during six daily activities performed by 24 participants. From this dataset, we selected three signal channels (G.x, G.y and G.z) from two actions (walk and jog) to construct a binary classification task, containing 80 samples with 200 time points each. The train/validation/test split is 56/8/16.
E.4 Introduction of Baselines